# Optimal Bound for PCA with Outliers using Higher-Degree Voronoi Diagrams

### v2

@@ 4 Randomized Algorithm
Using Lemma 5 and Lemma 4, the proof is immediate. As shown by Lemma 4, based on the definition of the Voronoi diagram (Definition 5), we have $\text{Ball}(L,\alpha)\subset VOD(L)$. Thus, $\mathbb{P}\left(VOD(L)\right)\geq\mathbb{P}\left(\text{Ball}(L,\alpha)\right)$, completing the claim. ∎
We conclude this section by presenting a theorem that highlights the effectiveness of our proposed randomized algorithm for solving PCA with Outliers problem. This algorithm leverages the intrinsic geometric properties of the Grassmannian manifold and the concept of the $\alpha$-Gap Constant to iteratively sample subspaces and identify outliers, ensuring a high probability of finding the optimal subspace. By analyzing the Voronoi cells associated with these subspaces, we construct a framework that conforms with the algorithm’s performance guarantees. The theorem shows that the algorithm can efficiently produce near-optimal solutions with high probability, making it a powerful tool for analyzing high-dimensional data.
@@
Let $X\subset\mathbb{R}^{d}$ be a set of vectors in a $d$-dimensional real vector space. There exists a randomized algorithm that, with probability at least $(1-\delta)^{T}$, produces an optimal solution for the PCA with Outliers problem. Here, $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$, where $\alpha$ is a constant, and $T$ is the number of subspaces sampled. The algorithm runs in time:
$2^{O(r(d-r))}\text{poly}(n,d),$
where $n$ is the number of vectors, $d$ is the dimensionality of the space, and $r$ is the target subspace dimensionality. The term $\text{poly}(n,d)$ accounts for the operations per sample, including PCA computation.
@@
Based on Theorem 3, we know that sampling multiple subspaces uniformly at random has a high probability of hitting the Voronoi cell that contains the optimal subspace. The steps of the algorithm are as follows:
Algorithm 2 Randomized PCA with Outliers
1:$\mathcal{T}\leftarrow$ uniformly sample $T$ subspaces of rank $r$
2:Compute the loss for each subspace in $\mathcal{T}$
3:$\hat{L}\leftarrow$ subspace with the minimal loss
4:$S\leftarrow$ set of $r$ vectors with the largest distances to $\hat{L}$
5:Perform PCA on the dataset $X\setminus S$ after removing the identified outliers
6:return Principal components of $X\setminus S$
Intuitively, by uniformly sampling subspaces, we increase the likelihood of selecting a subspace that is close to the optimal subspace. Each sampled subspace has an independent probability of $1-\delta$ of being within the optimal Voronoi cell. Therefore, by repeating this sampling process $T$ times, the probability that at least one sampled subspace lies within the optimal Voronoi cell is given by the complementary probability of all samples failing, which is $(1-\delta)^{T}$. To achieve a high probability of success (say, at least $1-\epsilon$), we want:
$(1-\delta)^{T}\geq 1-\epsilon.$
Taking the natural logarithm of both sides and using the approximation $\ln(1-x)\approx-x$ for small $x$:
$T(-\delta)\approx\ln(1-\epsilon),$
$T\approx\frac{\ln(1-\epsilon)}{-\delta}.$
Given that $\delta$ is small, $\ln(1-\epsilon)$ approximates $-\epsilon$, so we have $T\approx\frac{\epsilon}{\delta}$ which substituting $\delta=\mathbb{P}(\text{Ball}(L,\alpha))$:
$T\approx\frac{\epsilon}{\alpha^{r(d-r)}\left(\frac{2d}{r(d-r)}\right)^{r(d-r)/2}}.$
The above expression shows that $T$ is inversely proportional to the probability measure of the ball, which itself is a function of $\alpha$:
$T\propto\left(\frac{1}{\alpha}\right)^{r(d-r)}.$
As $\alpha$ decreases, the volume of the ball in the Grassmannian manifold also decreases exponentially. Hence, to maintain a high probability of hitting the optimal subspace, the number of samples $T$ must increase exponentially. We get:
$T\propto 2^{r(d-r)}.$
The computational complexity of this algorithm is $O(T)\times\text{poly}(n,d)$, where $T$ is the number of sampled subspaces, $n$ is the number of vectors, and $d$ is the dimensionality of the space. The polynomial term $\text{poly}(n,d)$ accounts for the operations per sample, such as computing the Principal Component Analysis (PCA) on each sampled subspace. This complexity arises from the repeated sampling and evaluation of subspaces, followed by the PCA computation on the reduced dataset. Given $T\propto 2^{r(d-r)}$, the complexity becomes:
$2^{O(r(d-r))}\text{poly}(n,d),$
highlighting that as the dimensionality reduction factor $\alpha$ decreases, the number of required samples $T$ increases exponentially, which in turn increases the overall computational burden. ∎
In summary, our randomized algorithm leverages the structure of the Grassmannian manifold and the properties of subspace projections to efficiently solve the PCA with Outliers problem with high probability. The approach combines theoretical guarantees with practical feasibility, making it a useful method for handling outliers in high-dimensional data.
@@ 5 Conclusion
In this paper, we introduced a new algorithm for the problem of PCA in the presence of outliers, leveraging the concept of higher-degree Voronoi diagrams to effectively detect outliers. Our approach builds on previous work by offering an exact solution with a complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$, which is more efficient than the solutions in [11] and [26]. Furthermore, our method provides a more intuitive framework to solve the PCA with Outliers problem by clearly delineating the roles of data partitioning and outlier detection. We also presented a randomized version of our algorithm, showing that achieving the optimal subspace can be possible through uniform sampling. While this randomized solution is currently less efficient than the one in [11], we suggest that adopting Gaussian sampling could enhance its accuracy and efficiency, as our main future study direction. Another direction is exploring the dual concept of Voronoi diagrams, specifically Delaunay triangulation, which could provide further insights into the PCA with Outliers problem. The constructive algorithm for Delaunay triangulation presented by [1] offers a promising avenue for investigation. Finally, we would like to link our work to other variants of PCA with outliers, namely online PCA [14, 15]. In particular, we are interested in Bandit PCA [20], where there are feedback loops for having better sequential projections, and where data samples are noisy (have outliers).

### v3

@@ 5 Randomized Algorithm
By Lemma 4, with high probability, $\text{dist}_{\min}\geq c_{1}\,n^{-1/(d-r)}.$ Set
$\epsilon=\frac{\alpha}{2}\,c_{1}\,n^{-1/(d-r)}\leq\frac{\alpha}{2}\,\text{dist}_{\min}.$
Then, by Lemma 3, if $\|P_{\hat{L}}-P_{L}\|\leq\epsilon$, the (nk)(n-k)-nearest neighbor set of $^{\hat{L}}$ is exactly $X\setminus S^{*}$. By Lemma 2, the probability that $\|P_{\hat{L}}-P_{L}\|\leq\epsilon$ is at least $c_{0}(d,r)\,\epsilon^{r(d-r)}.$ Substitute the bound for $\epsilon$:
$\begin{split}\Pr\Bigl[X\setminus S^{*}=(n-k)\text{-NN of }{\hat{L}}\Bigr]&\geq c_{0}(d,r)\left(\frac{\alpha}{2}\,c_{1}\,n^{-1/(d-r)}\right)^{r(d-r)}\\ &=c_{0}(d,r)\,\left(\frac{\alpha\,c_{1}}{2}\right)^{r(d-r)}n^{-r}.\end{split}$
Defining
$C(d,r,\alpha)=c_{0}(d,r)\,\left(\frac{\alpha\,c_{1}}{2}\right)^{r(d-r)},$
the theorem is proved. ∎
This theorem provides the theoretical foundation for the analysis of our randomized algorithm 2. To achieve a high-confidence guarantee, the algorithm performs multiple independent trials, each involving the random selection of a subspace $\hat{L}$ and the computation of distances to this subspace. By repeating this process many times, the algorithm ensures that the probability of at least one successful trial is at least $1-\delta$.
The time complexity of the algorithm is dominated by the number of trials $t$ and the cost of each trial, which includes sampling a random subspace, computing distances, and performing PCA.
Algorithm 2 Randomized PCA with Outliers
1:$X$ ($n$ vectors in $\mathbb{R}^{d}$), $t$ (# of trials), $k$ (# of outliers), $r$ (Goal dimension)
[…]
4:while $t\geq 0$ do
5:  $t\leftarrow t-1$.
6:  $\hat{L}\leftarrow$ uniformly sampled subspace of rank $r$.
7:  $\hat{S}\leftarrow$ the set of $k$ vectors from X with maximum distance to $\hat{L}$.
8:  if $\text{loss}(\text{PCA}(X\setminus\hat{S},r),X\setminus\hat{S})\leq\tilde{l}$ then
[…]
10:   $\tilde{S}\leftarrow\hat{S}$
11:  end if
12:end while
13:return $\text{PCA}(X\setminus\tilde{S},r)$
@@
Let $X\subset\mathbb{R}^{d}$ be a set of $n$ vectors, $r\leq d$, and $\delta\in(0,1)$ be given parameters. Then, the randomized algorithm 2 returns the optimal solution to the PCA with Outliers problem with probability at least $1-\delta$ in time
$O\Bigl(\frac{n^{r}\log(1/\delta)}{C(d,r,\alpha)}\text{poly}(d,r)\Bigr)$
where $\alpha$ is the same as in Lemma 1 and $C(d,r,\alpha)$ is an explicit constant (depending only on $d$, $r$, and $\alpha$) as in Theorem 3.
@@
By Theorem 3, a single random draw of an $r$-dimensional subspace lands in the optimal Voronoi cell with probability $p\geq\frac{C}{n^{r}}$. We perform $T$ independent trials of the sampling. Let the binary variable $X_{i}=1$ iff the $i$th trial succeeds. Then $S=\sum_{i=1}^{T}X_{i}$ has $\mathbb{E}[S]=Tp$. We want $\Pr[S\leq 0]\leq\delta$. By the Chernoff bound, for any $\epsilon\in(0,1)$,
$\Pr\bigl[S\leq(1-\epsilon)\mathbb{E}[S]\bigr]\leq\exp\Bigl(-\tfrac{\epsilon^{2}}{2}\mathbb{E}[S]\Bigr)\,.$
In particular, set $\epsilon=1$. Then $(1-\epsilon)\,\mathbb{E}[S]=0$, and $\Pr\bigl[S\leq 0\bigr]\leq\exp\bigl(-\tfrac{1}{2}\,Tp\bigr).$ Hence, choosing
$T\;=\;\Bigl\lceil\frac{2\,n^{r}\ln(1/\delta)}{C}\Bigr\rceil\;=\;O\!\bigl(n^{r}\log(1/\delta)/C\bigr)$
guarantees success with probability at least $1-\delta$.
Each trial requires $O\bigl(n\,\mathrm{poly}(d,r)\bigr)$ time, so the total running time is
$T\times O\bigl(n\,\mathrm{poly}(d,r)\bigr)\;=\;O\!\Bigl(\frac{n^{r}\log(1/\delta)}{C}\,\mathrm{poly}(d,r)\Bigr),$
Our analysis employs a conservative lower bound on the volume of the optimal Voronoi cell, which yields a cautious estimate for the algorithm’s success probability. In practice, the actual probability of sampling the optimal cell and the true success rate substantially exceeds this baseline bound.
@@ 6 Conclusion
In this paper, we introduced a new algorithm for the problem of PCA in the presence of outliers, leveraging the concept of higher-degree Voronoi diagrams to effectively detect outliers. Our approach builds on previous work by offering an exact solution with a complexity of $n^{d+\mathcal{O}(1)}\text{poly}(n,d)$, which is more efficient than the solutions in [13] and [27]. Furthermore, our method provides a more intuitive framework to solve the PCA with Outliers problem by clearly delineating the roles of data partitioning and outlier detection. We also presented a randomized version of our algorithm, showing that achieving the optimal subspace can be possible through uniform sampling. While this randomized solution is currently less efficient than the one in [13], we suggest that adopting Gaussian sampling could enhance its accuracy and efficiency, as our main future study direction. Our paper focused on theoretical aspect of the robust PCA, however we intend to implement our constructive algorithms for use in applications. Another direction is exploring the dual concept of Voronoi diagrams, specifically Delaunay triangulation, which could provide further insights into the PCA with Outliers problem. The constructive algorithm for Delaunay triangulation presented by [1] offers a promising avenue for investigation. Finally, we would like to link our work to other variants of PCA with outliers, namely online PCA [17, 18]. In particular, we are interested in Bandit PCA [22], where there are feedback loops for having better sequential projections, and where data samples are noisy (have outliers).